# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

### v1

@@ 5 Conclusions and Future Work
While the present study focused on multi-fidelity learning with just two or three quantum mechanical datasets, multi-fidelity learning is capable of leveraging several datasets with vastly different amounts of data. Future work should investigate alternative multi-fidelity learning training protocols that better account for data volume imbalance between different levels of theory. This study considered multi-fidelity learning when both the low-level and high-level data exist for the same molecular structures. While this approach clearly improves accuracy on the types of systems included in the training set, it is unlikely to improve transferability. Future work should consider how multi-fidelity learning can leverage disparate datasets to share knowledge from different regions of chemical space. Ideally, the molecular structures would cover completely different portions of the potential energy surface, although some overlap may be necessary for joint learning. [36] For example, future work will determine if multi-fidelity learning with low-level forces and high-level energies is beneficial when training to both non-reactive (e.g., ANI-1ccx) and reactive (e.g., Transition-1x) [10] datasets or when training to both in vacuo (e.g., ANI-1ccx) and periodic boundary (e.g., ANI-1xnr) [58] datasets. In light of the recent advances training with Hessian data, future work could also investigate multi-fidelity learning with low-level Hessians. [20, 21]
@@ Acknowledgments
All models in this study utilize the “Hierarchically Interacting Particle Neural Network with Tensor Sensitivity Information” (HIP-NN-TS) architecture, an extension of the original HIP-NN architecture. [38, 24] The original HIP-NN architecture is limited in terms of how information about atomic neighborhoods is collected by the model. The extended HIP-NN-TS architecture can incorporate either scalar, dipole, or quadrupole information about the feature distribution in the environment of an atom. [24] HIP-NN-TS with tensor sensitivity of zero (i.e., scalar information) is equivalent to the original HIP-NN model. The accuracy (and computational cost) of HIP-NN-TS increases somewhat with increasing tensor sensitivity. To maximize accuracy, all HIP-NN-TS models reported in the main text utilize tensor sensitivity of $\ell=2$ (that is, using scalar, dipole, and quadrupole information). For comparison, Supporting Information provides results for other HIP-NN-TS models with tensor sensitivity of zero (scalar information) and one (scalar and dipole information). For brevity, we refer to HIP-NN-TS simply as HIP-NN throughout the main text.
@@
where $E^{n}_{i}$ is the $n^{\rm th}$ hierarchical energy for atom $i$, $w$ is a trainable weight matrix and $b$ is a trainable bias vector. For $n=0$, $z_{i,a}$ corresponds to the one-hot encoding input feature vector. For $n>0$, $z_{i,a}$ corresponds to the transformed feature vector following the $n^{\rm th}$ interaction block. The hierarchical energies for a given atom $i$ are summed together to give the “fictitious” atomic energy

### v2

@@ 5 Conclusions and Future Work
While the present study focused on multi-fidelity learning with just two or three quantum mechanical datasets, multi-fidelity learning is capable of leveraging several datasets with vastly different amounts of data. Future work could investigate alternative multi-fidelity learning training protocols that better account for data volume imbalance between different levels of theory. This study considered multi-fidelity learning when both the low-level and high-level data exist for the same molecular structures. While this approach clearly improves accuracy on the types of systems included in the training set, it is unlikely to improve transferability. Future work could consider how multi-fidelity learning can leverage disparate datasets to share knowledge from different regions of chemical space. Ideally, the molecular structures would cover completely different portions of the potential energy surface, although some overlap may be necessary for joint learning. [50] For example, future work could determine if multi-fidelity learning with low-level forces and high-level energies is beneficial when training to both non-reactive (e.g., ANI-1ccx) and reactive (e.g., Transition1x) [18] datasets or when training to both in vacuo (e.g., ANI-1ccx) and periodic boundary (e.g., high-pressure MD) [80, 81, 82] datasets. Future work could also test the limits of multi-fidelity learning by using low-level forces that are computed with extremely inexpensive but even more approximate methods, such as classical force fields or semi-empirical quantum chemistry. [83] Future work could also consider including in the multi-fidelity dataset so-called “atomic energies.” Although atomic energies are not physical quantities that can be computed with quantum mechanical methods, previous studies have shown the advantage of (pre-)training to atomic energies [84, 85] computed with either a classical force field [86] or a “teacher” machine learning interatomic potential. [58, 87] Furthermore, in light of the recent advances training with Hessian data, [35, 36] future work could also investigate multi-fidelity learning with low-level Hessians.
@@ Acknowledgments
All models in this study utilize the “Hierarchically Interacting Particle Neural Network with Tensor Sensitivity Information” (HIP-NN-TS) architecture, [38] an extension of the original HIP-NN architecture. [52] The original HIP-NN architecture is limited in terms of how information about atomic neighborhoods is collected by the model. The extended HIP-NN-TS architecture can incorporate either scalar, dipole, or quadrupole information about the feature distribution in the environment of an atom. HIP-NN-TS with tensor sensitivity of zero (i.e., scalar information) is equivalent to the original HIP-NN model. The accuracy (and computational cost) of HIP-NN-TS increases somewhat with increasing tensor sensitivity. To maximize accuracy, all HIP-NN-TS models reported in the main text utilize tensor sensitivity of $\ell=2$ (that is, using scalar, dipole, and quadrupole information). For comparison, Supporting Information provides results for other HIP-NN-TS models with tensor sensitivity of zero (scalar information) and one (scalar and dipole information). For brevity, we refer to HIP-NN-TS simply as HIP-NN throughout the main text.
@@
where $E^{n}_{i}$ is the $n^{\rm th}$ hierarchical energy for atom $i$, $w$ is a trainable weight matrix and $b$ is a trainable bias vector. For $n=0$, $z_{i,a}$ corresponds to the one-hot encoding input feature vector. For $n>0$, $z_{i,a}$ corresponds to the transformed feature vector following the $n^{\rm th}$ interaction block. The hierarchical energies for a given atom $i$ are summed together to give the so-called atomic energy